how do you solve:
r + s + 2t - u = -3
2r + 3s + 3t + u = 2
4r + 2s - t + u = 5
s + 2t + 2u =7

Answer by Edwin McCravy(20086)

how do you solve:

 r +  s + 2t -  u = -3
2r + 3s + 3t +  u =  2
4r + 2s -  t +  u =  5
      s + 2t + 2u =  7

Get rid of the 2r by multiplying the 1st eq. 
temporarily by -2 and adding it to the 2nd eq.

-2[ r +  s + 2t -  u = -3
 1[2r + 3s + 3t +  u =  2
   4r + 2s -  t +  u =  5
         s + 2t + 2u =  7

 r +  s + 2t -  u = -3  
      s -  t + 3u =  8
4r + 2s -  t +  u =  5
      s + 2t + 2u =  7

Get rid of the 4r by multiplying the 1st eq.
temporarily by -4 and adding it to the 3rd eq.

-4[ r +  s + 2t -  u = -3  
         s -  t + 3u =  8
 1[4r + 2s -  t +  u =  5
         s + 2t + 2u =  7

    r +  s + 2t -  u = -3  
         s -  t + 3u =  8
       -2s - 9t + 5u = 17
         s + 2t + 2u =  7

Get rid of the -2s by multiplying the 2nd eq.
temporarily by 2 and adding it to the 3rd eq.

    r +  s + 2t -  u = -3  
     2[  s -  t + 3u =  8
     1[-2s - 9t + 5u = 17 
         s + 2t + 2u =  7

    r +  s + 2t -   u = -3  
         s -  t +  3u =  8
           -11t + 11u = 33 
         s + 2t +  2u =  7

Notice that the third equation can be divided
through by -11.  That will make things easier.

    r +  s + 2t -   u = -3  
         s -  t +  3u =  8
              t -   u = -3 
         s + 2t +  2u =  7

Get rid of the s in the bottom equation by 
multiplying the 2nd eq. temporarily by -1 and
adding it to the 4th eq.

    r +  s + 2t -  u = -3  
      -1[s -  t + 3u =  8
              t -  u = -3
       1[s + 2t + 2u =  7

    r +  s + 2t -  u = -3  
         s -  t + 3u =  8
              t -  u = -3
             3t -  u = -1

Getr rid of the 3t by multiplying the 3rd eq.
temporarily by -3 and adding it to the 4th eq.

    r +  s + 2t -  u = -3  
         s -  t + 3u =  8
          -3[ t -  u = -3
           1[3t -  u = -1

    r +  s + 2t -  u = -3  
         s -  t + 3u =  8
              t -  u = -3
                  2u =  8

Divide the 4th equation through by 2

    r +  s + 2t -  u = -3  
         s -  t + 3u =  8
              t -  u = -3
                   u =  4

The 4th equation tells us the value of u,
which is 4.  Now we're ready to do back-
substitution:

Substitute u = 4 into the 3rd
equation:

              t - u = -3
              t - 4 = -3
                  t =  1

Substitute t = 1 and u = 4 into the 2nd
equation:

         s -  t + 3u =   8
       s -  1 + 3(4) =   8
           s - 1 + 12 =  8
               s + 11 =  8
                    s = -3

Substitute s = -3, t = 1 and u = 4 into the 1st
equation:

     r +  s + 2t -  u = -3 
r + (-3) + 2(1) - (4) = -3
        r - 3 + 2 - 4 = -3
                r - 5 = -3
                    r =  2

So the solution is

(r, s, t, u) = (2, -3, 1, 4)  

Edwin